# RepoHyper: Better Context Retrieval Is All You Need for Repository-Level Code Completion

### v1

@@ Abstract
Code Large Language Models (CodeLLMs) have demonstrated impressive proficiency in code completion tasks. However, they often fall short of fully understanding the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies, which can result in less precise completions. To overcome these limitations, we present RepoHyper, a multifaceted framework designed to address the complex challenges associated with repository-level code completion. Central to RepoHyper is the Repo-level Semantic Graph (RSG), a novel semantic graph structure that encapsulates the vast context of code repositories. Furthermore, RepoHyper leverages Expand and Refine retrieval method, including a graph expansion and a link prediction algorithm applied to the RSG, enabling the effective retrieval and prioritization of relevant code snippets. Our evaluations show that RepoHyper markedly outperforms existing techniques in repository-level code completion, showcasing enhanced accuracy across various datasets when compared to several strong baselines.
@@
Huy N. Phan FPT Software AI Center, VinUniversity HuyPN16@fpt.com           Hoang N. Phan Department of Computer Science Nanyang Technological University C210055@e.ntu.edu.sg
  Tien N. Nguyen Computer Science Department The University of Texas at Dallas tien.n.nguyen@utdallas.edu          Nghi D. Q. Bui Fulbright University nghi.bui@fulbright.edu.vn
@@ 1 Introduction
To address these limitations, repository-level code completion has emerged as a promising solution. It aims to incorporate the full context of a project, including inter-file relationships, imported modules, and the overarching project structure Liu et al. (2023b); Liao et al. (2023); Shrivastava et al. (2023a); Agrawal et al. (2023); Shrivastava et al. (2023b). These methodologies generally employ a similarity-based approach to retrieve contexts for completing a given code snippet, drawing either from raw source code or a pre-constructed database with essential metadata. However, this strategy exhibits significant limitations. It often fails to consider that diverse contexts within the current repository, not necessarily involving similar code, can provide valuable insights for code completion. This includes the intricate network of dependencies, shared utility functions, inter-module method calls, class hierarchies, inter-class dependencies, and encapsulation patterns—all of which are fundamental to program semantics.
@@ 3.3 Re-ranking as Link Prediction
$\mathbf{C}=\text{ LLM }\left(Q,\left[\mathbf{P}_{g_{1}};\mathbf{P}_{g_{2}};\cdots;\mathbf{P}_{g_{N_{2}}}\right]\right)$ (4)
@@ 5.2 Performance on Code Completion
Gpt-3.5-turbo
In-File-Only 26.35 33.14 36.31 44.01
Gold-Only 30.59 38.37 40.65 49.12
@@
DeepSeek-Coder
In-File-Only 25.46 32.39 34.23 42.13
Gold-Only 27.29 35.15 37.10 46.08
@@ 6.1 Hyper-parameters Sensitivity and Effects
[…]
@@ 6.2 Retrieved Nodes Analysis
SimNParClsChiClsSibImpSibImpSimNImpParCls00.20.20.40.40.60.60.80.811Context TypesRetrieval Performance Acc@5RepoHyperSimilarity-based Search
@@ 7 Conclusion
Our evaluations, conducted on the RepoBench benchmark, demonstrates RepoHyper superior performance over state-of-the-art methods.

### v2

@@ Abstract
Code Large Language Models (CodeLLMs) have demonstrated impressive proficiency in code completion tasks. However, they often fall short of fully understanding the extensive context of a project repository, such as the intricacies of relevant files and class hierarchies, which can result in less precise completions. To overcome these limitations, we present RepoHyper, a multifaceted framework designed to address the complex challenges associated with repository-level code completion. Central to RepoHyper is the Repo-level Semantic Graph (RSG), a novel semantic graph structure that encapsulates the vast context of code repositories. Furthermore, RepoHyper leverages Expand and Refine retrieval method, including a graph expansion and a link prediction algorithm applied to the RSG, enabling the effective retrieval and prioritization of relevant code snippets. Our evaluations show that RepoHyper markedly outperforms existing techniques in repository-level code completion, showcasing enhanced accuracy across various datasets when compared to several strong baselines. Our implementation of RepoHyper can be found at https://github.com/FSoft-AI4Code/RepoHyper.
@@
Huy N. ${}^{\spadesuit}$  Hoang N. ${}^{\clubsuit}$  Tien N. ${}^{\diamondsuit}$ Nghi D. Q. ${}^{\heartsuit}$$^{\spadesuit}$ Software AI Center, Viet Nam ${}^{\clubsuit}$epartment of Computer Science, Nanyang Technological University, Singapore ${}^{\diamondsuit}$omputer Science Department, The University of Texas at Dallas, USA ${}^{\heartsuit}$ulbright University, Viet Nam huypn16@fpt.com, C210055@ntu.edu.sg, tien.n.nguyen@utdallas.edu, nghi.bui@fulbright.edu.vn
@@ 1 Introduction
To overcome these shortcomings, a direct but complex solution involves enhancing the context length of language models by applying efficient attention techniques Dao et al. (2022); Dao (2023); Press et al. (2021); Chen et al. (2023). Nonetheless, increasing the context length significantly raises costs and is not feasible indefinitely, especially with respect to the voluminous number of files in a given repository. Hence, there is a crucial need for more refined strategies that accurately identify relevant contexts rather than indiscriminately analyzing every file within a repository. In response to this challenge, the concept of repository-level code completion has gained traction. It aims to incorporate the full context of a project, including inter-file relationships, imported modules, and the overarching project structure Liu et al. (2023b); Liao et al. (2023); Shrivastava et al. (2023a); Agrawal et al. (2023); Shrivastava et al. (2023b). These methodologies generally employ a similarity-based approach to retrieve contexts for completing a given code snippet, drawing either from raw source code or a pre-constructed database with essential metadata. However, this strategy exhibits significant limitations. It often fails to consider that diverse contexts within the current repository, not necessarily involving similar code, can provide valuable insights for code completion. This includes the intricate network of dependencies, shared utility functions, inter-module method calls, class hierarchies, inter-class dependencies, and encapsulation patterns—all of which are fundamental to program semantics.
@@ 3.3 Re-ranking as Link Prediction
[…]
@@ 5.2 Performance on Code Completion
Gpt-3.5-turbo ${}^{\ast}$ 26.35 33.14 36.31 44.01
${}^{\ast}$ 30.59 38.37 40.65 49.12
@@
DeepSeek-Coder ${}^{\ast}$ 25.46 32.39 34.23 42.13
${}^{\ast}$ 27.29 35.15 37.10 46.08
@@ 6.1 Hyper-parameters Sensitivity and Effects
kNN D=M=1, K=0.35*|$\mathcal{G}$| 53% 35%
Table 3: Sensitivity Analysis. $D$ : maximum depth; $M$ : maximum number of nodes; $K$ : in $\mathrm{kNN}$ algorithm. During expansion, the search algorithm will expand $K$ nodes to Coverage % of the number nodes of the repolevel semantic graph, then extract explored node into a sub-graph. There’s a Hits % probability that there’s an optimal context in this sub-graph. In kNN baseline, we use $K$=$35\%$ of size of the original graph.
@@ 6.2 Retrieved Nodes Analysis
ptSimNptParClsptChiClsptSibptImpSibptImpSimNptImpParCls000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Context TypesRetrieval Performance Acc@5RepoHyperSimilarity-based Search
@@ 7 Conclusion
Our evaluations, conducted on the RepoBench benchmark, demonstrate RepoHyper’s superior performance over state-of-the-art methods.